# Theoretical Insights into Line Graph Transformation on Graph Learning

### v1

@@ Abstract
Graph isomorphism testing, Strongly regular graphs, CFI graphs, Expressivity, Graph neural networks.
editors: List of editors’ names
@@ 1 Introduction
Graph Neural Networks (GNNs) are prominent models widely studied in the domain of graph representation learning, with various applications ranging from social network analysis (Fan et al., 2019) and recommendation systems (Baskin et al., 2016) to drug discovery (Muzio et al., 2020). In particular, advancements in graph theory have played a significant role in the evolution of GNNs. The integration of graph measures, such as graph spectrum (Baldesi et al., 2018), centrality (Maurya et al., 2021), and modularity (Tsitsulin et al., 2023), has played a crucial role in shaping the design, analysis, and theoretical foundations of GNN research.
@@
In this study, we aim to understand the effects of line graph transformations on the data from a graph theory perspective. We first observe that conducting a graph isomorphism test between a pair of connected graphs is equivalent to conducting it between their respective line graphs with the exception of the pair of graphs isomorphic to $K_{1,3}$ and $C_{3}$. Note that this holds for repeated applications of line graph transformation. These observations inspire us to demonstrate the equivalence of the graph isomorphism test and uniform function approximation on graphs after an arbitrary number of transformations $L^{(n)}(G)$ and the original graph $G$.
Furthermore, we shift our focus to Cai-Fürer-Immerman (CFI) graphs and strongly regular graphs, which are notably difficult for the WL test. We demonstrate that with several applications of line graph transformation, the resulting line graph will no longer exhibit the same property, potentially simplifying the process of differentiation through standard isomorphism tests or GNNs. We then perform a series of experiments to evaluate how line graph transformations affect the performance of standard isomorphism tests and GNNs in distinguishing difficult graph classes, focusing on the effectiveness of applying lower-order WL tests on the transformed line graphs for isomorphism checks.
@@
$a$$b$$c$$d$$G$$a$$b$$c$$d$$ab$$ac$$bc$$bd$$ac$$ab$$bc$$bd$$L(G)$
Figure 1: An example of converting a graph $G$ to its line graph $L(G)$.
@@ 6 Conclusions
The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558
@@ F.1 Code availability
We open-sources our code to replicate the experiments on GitHub. The details about the experimental setup and training parameters can be found in the GitHub repository.

### v2

@@ Abstract
Graph isomorphism testing, Strongly regular graphs, CFI graphs, Expressivity, Graph neural networks, Line graphs.
editors: Simone Azeglio, Christian Shewmake, Bahareh Tolooshams, Sophia Sanborn, Chase van de Geijin, Nina Miolane
@@ 1 Introduction
Graph Neural Networks (GNNs) are prominent models widely studied in the domain of graph representation learning, with various applications ranging from social network analysis (Fan et al., 2019) and recommendation systems (Baskin et al., 2016) to drug discovery (Muzio et al., 2020). In particular, advancements in graph theory have played a significant role in the evolution of GNNs. The integration of graph measures, such as graph spectrum (Baldesi et al., 2018), centrality (Maurya et al., 2021), and modularity (Tsitsulin et al., 2023), has shaped the design, analysis, and theoretical foundations of GNN research.
@@
$a$$b$$c$$d$$G$$a$$b$$c$$d$$ab$$ac$$bc$$bd$$ac$$ab$$bc$$bd$$L(G)$
Figure 1: An example of converting a graph $G$ to its line graph $L(G)$.
@@
In this study, we aim to understand the effects of line graph transformations on the data from a graph theory perspective. We first observe that conducting a graph isomorphism test between a pair of connected graphs is equivalent to conducting it between their respective line graphs with the exception of the pair of graphs isomorphic to $K_{1,3}$ and $C_{3}$. Moreover, this holds for repeated applications of line graph transformation. These observations inspire us to demonstrate the equivalence of the graph isomorphism test and uniform function approximation on graphs after an arbitrary number of transformations $L^{(n)}(G)$ and the original graph $G$.
Furthermore, we shift our focus to Cai-Fürer-Immerman (CFI) graphs and strongly regular graphs, which are notably difficult for the WL test. We demonstrate that with several applications of line graph transformation, the resulting line graph no longer exhibits the same property, potentially simplifying the process of differentiation through standard isomorphism tests or GNNs. We then perform a series of experiments to evaluate how line graph transformations affect the performance of standard isomorphism tests and GNNs in distinguishing difficult graph classes, focusing on the effectiveness of applying lower-order WL tests on the transformed line graphs for isomorphism checks.
@@ 6 Conclusions
The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558. In addition, F.Y. would like to thank Dr. Seth Flaxman for his support in attending the conference.
@@ F.1 Code availability
We open-source our code to replicate the experiments on GitHub. The details about the experimental setup and training parameters can be found in the GitHub repository.